OpenStudy (ttop0816):

Simplify the following expression. sec^2x + sin^2x - tan^2x + cos^2x -2 0 1 2

hartnn (hartnn):

just pythagorean identities \(\Large \sin^2x+\cos^2x =1 \\ \Large \sec^2x -\tan^2x =1\)

hartnn (hartnn):

sec^2x + sin^2x - tan^2x + cos^2x = (sec^2x - tan^2x) + (sin^2x + cos^2x) =...

OpenStudy (ttop0816):

@hartnn 2??

hartnn (hartnn):

absolutely :)